OpenStudy (anonymous):

Solve 3x - 4 ≤ 2 or 2x + 11 ≥ -1

OpenStudy (anonymous):

3x - 4 <= 2 3x <= 2 + 4 3x <= 6 x <= 2 2x + 11 >= -1 2x >= -1 - 11 2x >= - 12 x >= -6